OpenStudy (anonymous):

Square root of 3d-8=d-2

OpenStudy (anonymous):

Are you sure that's the right equation? √(3d-8=d-2) ?

OpenStudy (anonymous):

YES .

OpenStudy (anonymous):

I don't know then for this one. :( I'm sorry.

OpenStudy (anonymous):

Well Lyssah Lets Solve this together

OpenStudy (akshay_budhkar):

even i will help ishaan.. lol.. hi ishaan

OpenStudy (anonymous):

I guess we could give it a go .

OpenStudy (akshay_budhkar):

√(3d-8=d-2)?????????? or \[\sqrt{3d-8}=d-2\]

OpenStudy (anonymous):

^ That's what I wanted to know.

OpenStudy (akshay_budhkar):

it cant be the first one lyssah.. please confirm the question:)

OpenStudy (akshay_budhkar):

what are you typing ishaan???

OpenStudy (anonymous):

- Yes , that .

OpenStudy (anonymous):

Ok Is it \[\sqrt(3d -8) = \sqrt (d-2)\]

OpenStudy (anonymous):

Damn Latex is not working

OpenStudy (anonymous):

Haha Worked

OpenStudy (akshay_budhkar):

ishaan that cant be the question.. as far as i know on what the person has asked in next question it is the 2nd one

OpenStudy (anonymous):

=d-2 shouldnt be undr a bracket thingy .

OpenStudy (anonymous):

ok Lyssah

OpenStudy (anonymous):

Square then Both sides

OpenStudy (anonymous):

Im sooo confused .